In a stream of glycerine in motion, at a certain point the velocity gradient is 0.25 metre per sec per metre . The mass density of fluid is 1268.4 kg per cubic metre and kinematic viscosity is 6.30 x 10^-4 square metre per second. Calculate the shear stress at the point.
1
Expert's answer
2020-11-09T20:29:19-0500

By the Newton's Law of Viscosity

"\tau=\mu\dfrac{du}{dy}"

Kinematic viscosity is defined as the ratio between the dynamic viscosity and and density of fluid.


"\nu=\dfrac{\mu}{\rho}"

Substitute


"\tau=\nu\cdot\rho\cdot\dfrac{du}{dy}"

Given

"\rho=1268.4kg/m^3,"

"\nu=6.3\times10^{-4}m^2/s ,"

"\dfrac{du}{dy}=0.25/s"

"\tau=1268.4kg/m^3\cdot6.3\times10^{-4}m^2/s\cdot0.25/s=""=0.199773N/m^2\approx0.2N/m^2"

The shear stress at the point is "0.2N\/m^2."